To split (s+10)/((s+4)(s-2)) into two terms.
1
Expert's answer
2020-05-06T18:44:17-0400

"\frac{s+10}{(s+4)(s-2)} = \frac{A}{s+4}+\frac{B}{s-2}"


"\frac{s+10}{(s+4)(s-2)} = \frac{A(s-2)+B(s+4)}{(s+4)(s-2)}"


"s+10 = As-2A+Bs+4B"


"A + B = 1"

"4B-2A=10"


"A= 1-B"

"4B-2(1-B)=10"


"4B-2+2B=10"

"6B=12"

"B=2"


"A=1-2=-1"


Answer

"\frac{-1}{s+4}+\frac{2}{s-2}"